OpenStudy (anonymous):

Help please Which fraction is equivalent to 4 2/3? is it 14/3

OpenStudy (butterfield1215):

yep

OpenStudy (anonymous):

1 whole number is 3/3 so

OpenStudy (anonymous):

yeh

OpenStudy (anonymous):

ok but thats not the answer on the quiz

OpenStudy (butterfield1215):

3*4 is12+2 is14

OpenStudy (butterfield1215):

so 14/3

OpenStudy (anonymous):

Which fraction is equivalent to 4 2/3?

OpenStudy (butterfield1215):

14/3 silly

OpenStudy (anonymous):

sorry trying to get the answers

OpenStudy (anonymous):

A. 13/12 B. 26/12 C. 13/6 D. 28/6

OpenStudy (butterfield1215):

oh ok

OpenStudy (anonymous):

those are what i see lol theres no that

OpenStudy (butterfield1215):

d

OpenStudy (anonymous):

d i believe cuz 3 x 2 is 6 and 14 x 2 is 28

OpenStudy (anonymous):

so you get 28/6

OpenStudy (vianne):

D

OpenStudy (vianne):

Some equivalent fractions of 14/3 are: 14/3 = 28/6 = 42/9 = 56/12 = 70/15 = 84/18 = 98/21 = 112/24 = 126/27 = 140/30 = 154/33 = 168/36 = 182/39 = 196/42 = 210/45 = 224/48 = 238/51 = 252/54 = 266/57 = 280/60

OpenStudy (anonymous):

ok thanks i hope your not lying

OpenStudy (vianne):

Nope, I even backed up my problem.. I did paper + online. I used online calculator + did it on paper. So it's 100% right I think!